f(X)=2x+1, if x,2
=3x-2,if x>or=2 then f(f(X))=

Grade:12th pass

1 Answers

Sher Mohammad IIT Delhi
10 years ago
f(x)=2x+1 x<2
=3x-2 otherwiese
f(0)=1
f(1)=3
f(2)= 4
f(3)=7
f(f(x))= 2f(x)+1 f(x)<2 imply x<=0
=4x+5 x<=0
=3f(x)-2 f(x)>=2 imply x>0
=9x-8 x>0

sher mohammad
b.tech, iit delhi